OpenStudy (ttop0816):

Solve the following system of equations graphically. x + y - 4 = 0 x - y = 0 The solution lies in quadrant _____. I II III IV

Directrix (directrix):

Are you required to draw the graphs?

Directrix (directrix):

If not, you can look at the two equations and "see" the solution. x + y =4 x - y = 0 -------- To numbers that add to 4 and subtract to 0. The question is, of course, in which quadrant is the solution.

OpenStudy (ttop0816):

its quad 4 then??

OpenStudy (anonymous):

Well, if the question says graphically, you may want to sketch the graphs of these 2 equations but you can also analyze the two equations. Convert them to y=4-x and y=x, then you know that the second equation passes through the i and iii quadrants, and then the first equation passes through i, ii and iii quadrants as you can see when x=0 y=4. then the intersect must be in the i quadrant... I hope I'm right, I've zoomed out many times today~~

OpenStudy (ttop0816):

@Takuno so your saying its in Quadrant 1??

OpenStudy (anonymous):

Yes, if it's asking for the intersect. Because when you have x+y=o indicates that y=-x, the graph of this equation is in quadrant ii and iv, but the equation says x+y-4=o then y=4-x which is equal to the graph of y=x goes up for 4 unit which shows that it passes through quadrant i, ii and iv(sorry that I typied qiii above, it should be i, ii and iv).

OpenStudy (anonymous):

Plot the graphs and you will see it.

OpenStudy (anonymous):

@Miyuru incisive O(∩_∩)O

OpenStudy (anonymous):

?